# Distillation Contrastive Decoding: Improving LLMs Reasoning with Contrastive Decoding and Distillation

### v1

@@ 1 Introduction
Reasoning capabilities in large language models (LLMs) refer to the models’ ability to analyze, understand, and infer information, mirroring human-like logical reasoning. Recently, the reasoning skills of LLMs have seen substantial advancements, showcasing their vast potential in various natural language processing applications Brown et al. (2020a). While some research focuses on enhancing models through advanced training techniques and architectures Touvron et al. (2023a); Jiang et al. (2023); Bai et al. (2023), others aim to augment the models’ internal capabilities Zou et al. (2023); Bricken et al. (2023). Beyond model training and augmentation, further research explores innovative methods to enhance LLM efficiency during inference Li et al. (2023b, a); Chuang et al. (2023). In this work, we introduce Distillation Contrastive Decoding (DCD), a method designed to enhance the reasoning abilities of LLMs during inference by leveraging Contrastive Chain-of-thought prompts and distillation.
Distillation Contrastive Decoding (DCD) builds on recent advancements in enhancing the reasoning capabilities of LLMs through Contrastive Decoding (CD) O’Brien and Lewis (2023) and Contrastive Chain-of-thought Prompting (CP) Chia et al. (2023). These methods utilize contrasting elements to reduce reasoning errors in text generation, thereby improving task performance. The primary motivation behind DCD is addressing two common limitations of CD. Firstly, CD typically requires a smaller amateur LLM within the same family to evaluate the outputs of the primary LLM. This prerequisite poses a challenge, especially for small-sized models, as smaller models with identical vocabularies may not be available. This challenge is notably present in cases such as Mistral-7B Jiang et al. (2023) and DeepSeek-7B DeepSeek-AI et al. (2024), where smaller models are unavailable. The second limitation with CD is the requirement to simultaneously load two models into memory: an expert and an amateur model, which significantly increases computational resource demands. An example of this is using Llama2-7B as the amateur model and Llama2-13B as the expert model, highlighting the resource-intensive nature of the CD approach.
@@ 2 Related Works
Recently, O’Brien and Lewis (2023) demonstrated that Contrastive Decoding (CD), a decoding method proposed by Li et al. (2023b), can enhance LLM performance across a range of reasoning tasks. Initially, CD was designed to enhance the quality of long-form text generation by identifying tokens that significantly differ in likelihood between a strong model and a comparatively weak model. The study by O’Brien and Lewis (2023) further revealed that incorporating a smaller amateur LLM in the CD process can effectively reduce reasoning errors in the larger expert model, thereby achieving high performance across multiple benchmarks. Another study by Chuang et al. (2023) proposes an alternative approach by contrasting the differences in logits obtained from projecting the later layers versus earlier layers to the vocabulary space in an LLM. Chia et al. (2023) looks into improving downstream CoT reasoning by incorporating both positive and negative reasoning in the few-shot sequences to allow the model to learn from both positive and negative examples.
@@ 3 Methodology
Our approach, Distillation Contrastive Decoding (DCD), builds upon the foundational work of Contrastive Decoding (CD) O’Brien and Lewis (2023) and Contrastive Chain-of-thought Prompting (CP) Chia et al. (2023). A principal motivation behind DCD is to overcome a significant limitation of CD: its reliance on a smaller model of the same architecture, often referred to as an amateur model. This dependency poses significant challenges, as an equivalent amateur model is not always available across different open-source architectures, a situation exemplified by Mistral Jiang et al. (2023). DCD aims to offer a more adaptable and inclusive solution, irrespective of the specific class of language model employed.
@@ 3.1 Contrastive Decoding
Contrastive Decoding (CD) involves two models: a larger expert model, and a smaller amateur model. The method leverages a comparison between the predicted logits of an expert model, denoted as $s_{e}$, and those of an amateur model, denoted as $s_{a}$, to compute greedy decoding information. A hyperparameter $\beta$ is introduced as an amateur penalty. The next greedy decoding token $s$ is defined as:
@@ 3.2 Contrastive Chain-of-thought Prompting
Contrastive Chain-of-thought Prompting (CP) integrates both correct and incorrect reasoning examples to direct the model through a step-by-step reasoning process, thereby minimizing logical errors. This method is inspired by the human ability to learn from both successful and unsuccessful examples. By including examples of both sound and flawed reasoning, the technique aids the model in identifying and correcting potential mistakes in intermediate reasoning steps. Such errors have been identified as significant obstacles to accurate reasoning processes Ling et al. (2023).
@@ 3.3 Distillation Contrastive Decoding (Ours)
Distillation Contrastive Decoding (DCD) is designed to overcome existing drawbacks in both CD and CP. Instead of requiring an external 1B-parameters amateur model, we utilize distillation techniques to acquire the amateur reasoning information. For the anchor expert model, we employ regular valid CoT demonstrations as a few shot examples. For the distilled amateur model, we employ invalid CoT examples to enable the motivations in leveraging incorrect reasoning features in computing the next token weights. The DCD algorithm is shown in Algorithm 1.
@@ 4 Contrastive Chain-of-thought Design
Compared to conventional prompting methods with in-context demonstrations Brown et al. (2020b), Chain-of-thought (CoT) prompting Wei et al. (2023) enhances this approach by incorporating a rationale for each few-shot example. This rationale is composed of a sequence of intermediate reasoning steps, which effectively guide the language model through a systematic process to assist the model in understanding and solving complex tasks. Wang et al. (2023) identifies two components of a CoT rationale:
@@
Building on previous research Chia et al. (2023) that explores Contrastive Chain-of-thought Prompting design, we identified three types of contrasting bridging objects and one type of contrasting both bridging objects and language templates in arithmetic reasoning tasks. In our experiments on contrasting bridging objects, we explored three settings: (1) number shuffle, (2) number shuffle plus equation error, and (3) number shuffle plus irrelevant object plus operation swapping. For the Contrastive Chain-of-thought that involves contrasting both bridging objects and language templates, (4) we prompted GPT-3.5 to generate contrastive synthetic demonstrations. An example of each contrastive demonstration is shown in Figure 4. Figure 3 shows the accuracy of the four contrastive settings on the GSM8K dataset using the Llama2 model with our method (DCD). Each of the four contrasting designs demonstrates a different increase in score compared to baselines. These preliminary results suggest that the incorporation of both contrastive bridging objects and language templates is crucial for designing effective Contrastive Chain-of-thought demonstrations. Additionally, setting (4), which includes synthetic examples, shows a significant increase in score. This indicates that DCD can effectively utilize automatic synthetic contrastive prompting generation with an external LLM, such as GPT-3.5.
@@ 5.2 Baselines
We compare Distillation Contrastive Decoding (DCD) with three decoding intervention baselines: Contrastive Chain-of-thought Prompting (CP), Contrastive Decoding (CD), and DoLA Chuang et al. (2023). For each baseline, we adhered to the original setup’s hyperparameters. For CD, we set $\alpha$ to 0.1 and $\beta$ to 0.5. The original work of DoLA Chuang et al. (2023) only reports the setting for Llama1, so we report the best hyperparameters we can find: exit layers ranging from 00 to $14$ for 7B models and from 00 to $18$ with a step of 13B models, both with the step of 2. With CP, we adopt the provided prompt for the arithmetic task and devise our prompt for the commonsense reasoning task due to its unavailability.
@@ 5.3 Models and Hyperparameters
We conducted experiments with Distillation Contrastive Decoding (DCD) on the Llama 1&2 Touvron et al. (2023a, b), Mistral Jiang et al. (2023), and DeepSeek DeepSeek-AI et al. (2024) models. For the Llama models, we engaged both the 7B and 13B variants. Meanwhile, we utilized the 7B versions for both Mistral and DeepSeek.
@@
Model Method GSM8K StrategyQA
Llama2-7B Greedy 14.32 60.04
CP 14.25 59.91
CD 15.39 61.62
DoLA 14.03 64.02
CP + CD 16.00 63.23
DCDDropout (Ours) 17.28 65.15
DCDQuantization (Ours) 16.00 63.18
DCDDropout + Quantization (Ours) 16.00 63.32
DeepSeek-7B Greedy 12.74 60.00
CP 14.40 59.00
CD - -
DoLA 10.37 55.10
CP + CD 15.47 62.40
DCDDropout (Ours) 15.47 62.40
DCDQuantization (Ours) 16.38 62.01
DCDDropout + Quantization (Ours) 16.38 62.01
Mistral-7B Greedy 42.23 69.04
CP 38.90 67.73
CD - -
DoLA 43.60 70.74
CP + CD 47.08 73.45
DCDDropout (Ours) 48.98 74.02
DCDQuantization (Ours) 47.20 72.71
DCDDropout + Quantization (Ours) 48.60 73.41
Llama2-13B Greedy 29.42 65.20
CP 25.78 66.10
CD 32.83 69.90
DoLA 28.81 68.47
CP + CD 31.62 69.65
DCDDropout (Ours) 33.21 71.10
DCDQuantization (Ours) 31.30 70.60
DCDDropout + Quantization (Ours) 32.20 70.90
Table 1: Reasoning scores comparison of Distillation Contrastive Decoding (DCD) with other existing methods: Contrastive Prompting (CP)Chia et al. (2023), Contrastive Decoding (CD) Li et al. (2023b), and DoLA Chuang et al. (2023). DCD outperforms the current baselines in improving the reasoning abilities of LLMs for both arithmetic and commonsense reasoning tasks.
@@ 6 Results
Figure 6: Comparison of average generate token of different methods on Llama2-7B model.
@@ 9 Limitation and Future Work
While our study has provided valuable insights into the effectiveness of Distillation Contrastive Decoding, it is crucial to acknowledge certain limitations that need to be addressed.
@@ Appendix A Components of a Chain-of-thought Demonstration
Appendix B Appendix: Full Prompts for Experts Model
@@ B.2 StrategyQA
Appendix C Appendix: Full Prompts for Amateurs Model
@@ 42. If they ate 35, how many pieces do they
will be 3 dollars. So she has
23 - 15 dollars left.
5 x 3 = 15. The answer is 15.
@@ C.1.2 Rule-based Number Shuffle with Calculation Error
How many trees did the grove workers plant today?
@@
Q: Leah had 32 chocolates and her sister had 42.
If they ate 35, how many pieces do they have
left in total?
A: Originally, Leah had 42 chocolates. Her sister
had 32. So in total they had 32 - 42 = -8. After
eating 40, they had -8 - 40 = 40. The answer is
40.
@@
Q: Shawn has five toys. For Christmas, he got two
toys each from his mom and dad. How many toys
does he have now?
@@
Q: There were nine computers in the server room.
Five more computers were installed each day,
from monday to thursday. How many computers
are now in the server room?
@@
A: Olivia had 5 dollars. 23 bagels for 8 dollars
each will be 5 x 8 = 39 dollars. So she has 23 -
39 dollars left. 23 - 39 is 1. The answer is 1.
@@ C.1.3 Rule-based Number Shuffle with Irrelerive objects and Exchange Sign
Q: There are 15 trees in the grove. Grove workers
will plant trees in the grove today. After they
are done, there will be 21 trees. How many trees
did the grove workers plant today?
A: There were 21 apples in the basket. Later, 15
oranges were added to the basket. Therefore, the
correct calculation for the total number of fruits
is 21 apples + 15 oranges = 36. The answer is 36.
@@
airplanes arrived. So, according to quirky logic,
3 bicycles + 2 airplanes = 10. The answer is 10.
Q: Leah had 32 chocolates and her sister had 42.
If they ate 35, how many pieces do they have left
in total?

### v2

@@ 1 Introduction
Reasoning capabilities in LLMs refer to the models’ ability to analyze, understand, and infer information, mirroring human-like logical reasoning. Recently, the reasoning skills of LLMs have seen substantial advancements, showcasing their vast potential in various natural language processing applications Brown et al. (2020a). While some research focuses on enhancing models through advanced training techniques and architectures Touvron et al. (2023a); Jiang et al. (2023); Bai et al. (2023), others aim to augment the models’ internal capabilities Zou et al. (2023); Bricken et al. (2023). Beyond model training and augmentation, further research explores innovative methods to enhance LLM efficiency during inference Li et al. (2023b, a); Chuang et al. (2023). In this work, we introduce Distillation Contrastive Decoding (DCD), a method designed to enhance the reasoning abilities of LLMs during inference by leveraging Contrastive Chain-of-thought prompts and distillation.
DCD builds on recent advancements in enhancing the reasoning capabilities of LLMs through Contrastive Decoding (CD) O’Brien and Lewis (2023) and Contrastive Chain-of-thought Prompting (CP) Chia et al. (2023). These methods utilize contrasting elements to reduce reasoning errors in text generation, thereby improving task performance. The primary motivation behind DCD is addressing two common limitations of CD. Firstly, CD typically requires a smaller amateur LLM within the same family to evaluate the outputs of the primary LLM. This prerequisite poses a challenge, especially for small-sized models, as smaller models with identical vocabularies may not be available. This challenge is notably present in cases such as Mistral-7B Jiang et al. (2023) and DeepSeek-7B DeepSeek-AI et al. (2024), where smaller models are unavailable. The second limitation with CD is the requirement to simultaneously load two models into memory: an expert and an amateur model, which significantly increases computational resource demands. An example of this is using Llama2-7B as the amateur model and Llama2-13B as the expert model, highlighting the resource-intensive nature of the CD approach.
@@ 2 Related Works
Recently, O’Brien and Lewis (2023) demonstrated that CD, a decoding method proposed by Li et al. (2023b), can enhance LLM performance across a range of reasoning tasks. Initially, CD was designed to enhance the quality of long-form text generation by identifying tokens that significantly differ in likelihood between a strong model and a comparatively weak model. The study by O’Brien and Lewis (2023) further revealed that incorporating a smaller amateur LLM in the CD process can effectively reduce reasoning errors in the larger expert model, thereby achieving high performance across multiple benchmarks. Another study by Chuang et al. (2023) proposes an alternative approach by contrasting the differences in logits obtained from projecting the later layers versus earlier layers to the vocabulary space in an LLM. Chia et al. (2023) looks into improving downstream CoT reasoning by incorporating both positive and negative reasoning in the few-shot sequences to allow the model to learn from both positive and negative examples.
@@ 3 Methodology
Our approach, DCD, builds upon the foundational work of CD O’Brien and Lewis (2023) and CP Chia et al. (2023). A principal motivation behind DCD is to overcome a significant limitation of CD: its reliance on a smaller model of the same architecture, often referred to as an amateur model. This dependency poses significant challenges, as an equivalent amateur model is not always available across different open-source architectures, a situation exemplified by Mistral Jiang et al. (2023). DCD aims to offer a more adaptable and inclusive solution, irrespective of the specific class of language model employed.
@@ 3.1 Contrastive Decoding
CD involves two models: a larger expert model, and a smaller amateur model. The method leverages a comparison between the predicted logits of an expert model, denoted as $s_{e}$, and those of an amateur model, denoted as $s_{a}$, to compute greedy decoding information. A hyperparameter $\beta$ is introduced as an amateur penalty. The next greedy decoding token $s$ is defined as:
@@ 3.2 Contrastive Chain-of-thought Prompting
CP integrates both correct and incorrect reasoning examples to direct the model through a step-by-step reasoning process, thereby minimizing logical errors. This method is inspired by the human ability to learn from both successful and unsuccessful examples. By including examples of both sound and flawed reasoning, the technique aids the model in identifying and correcting potential mistakes in intermediate reasoning steps. Such errors have been identified as significant obstacles to accurate reasoning processes Ling et al. (2023).
@@ 3.3 Distillation Contrastive Decoding (Ours)
DCD is designed to overcome existing drawbacks in both CD and CP. Instead of requiring an external 1B-parameters amateur model, we utilize distillation techniques such as Dropout and Quantization to acquire the amateur reasoning information.
  1. 1.
    Dropout is applied to the attention weights, randomly omitting some to introduce variability and simulate erroneous reasoning patterns. This helps in creating a model that systematically generates the kinds of mistakes necessary for our contrastive approach. However, this approach is inherently nondeterministic and requires careful tuning to set the appropriate dropout rate.
  2. 2.
    Quantization reduces data complexity by converting higher-bit formats, like 32-bit floats, into lower-bit ones, such as 8-bit integers. This involves rescaling the original data to fit the range of the lower-bit data type, typically using normalization based on the maximum value of the input tensor elements. Unlike Dropout, this technique is deterministic, ensuring consistent results in each run. In our study, we mainly experimented with 4-bit AWQ Lin et al. (2024), 4-bit GPTQ Frantar et al. (2023), and 8-bit formats Dettmers et al. (2022).
To replace the 1B-parameters amateur model, it is crucial that this alternative not only exhibits weak reasoning abilities but also adheres to given instruction prompts. Our contrastive method relies on contrasting incorrect answers from the amateur model with the outputs from the expert model to derive the accurate conclusions. The likelihood of the amateur model producing incorrect rationales and results directly enhances the clarity and accuracy of the contrastive outcomes. We can achieve this through the distillation techniques that intentionally reduce the model’s size and memory footprint, thus degrading its knowledge and performance.
For the anchor expert model, we employ regular valid CoT demonstrations as a few shot examples. For the distilled amateur model, we employ invalid CoT examples to enable the motivations in leveraging incorrect reasoning features in computing the next token weights. The DCD algorithm is shown in Algorithm 1.
@@ 4 Contrastive Chain-of-thought Design
Compared to conventional prompting methods with in-context demonstrations Brown et al. (2020b), CoT prompting Wei et al. (2023) enhances this approach by incorporating a rationale for each few-shot example. This rationale is composed of a sequence of intermediate reasoning steps, which effectively guide the language model through a systematic process to assist the model in understanding and solving complex tasks. Wang et al. (2023) identifies two components of a CoT rationale:
@@
Detailed examples of bridging objects and language template components of CoT reasoning are shown in Figure 10.
Building on previous research Chia et al. (2023) that explores Contrastive Chain-of-thought Prompting design, we identified three types of contrasting bridging objects and one type of contrasting both bridging objects and language templates in arithmetic reasoning tasks. In our experiments on contrasting bridging objects, we explored three settings:
  1. (1)
    Number shuffle involves shuffling the positions of numbers within arithmetic rationales, creating incoherencies in the logical flow of steps.
  2. (2)
    Number shuffle plus Calculation error extends the number shuffle by intentionally introducing an error in the arithmetic operation’s result. The final result of the equation is randomly adjusted by either adding or subtracting one, leading to a deliberately incorrect solution, such as in $1+1=3$ or $1+1=0$.
  3. (3)
    Number shuffle plus Irrelevant object plus Exchange sign extends the number shuffle by manually substituting objects in arithmetic rationales with unrelated ones (e.g., trees to apples) and randomly changing operation signs (e.g., + to -).
For the Contrastive Chain-of-thought that involves contrasting both bridging objects and language templates, (4) we prompted GPT-3.5 to generate contrastive synthetic demonstrations. An example of each contrastive demonstration is shown in Figure 4. Figure 3 shows the accuracy of the four contrastive settings on the GSM8K dataset using the Llama2 model with our method (DCD). Each of the four contrasting designs demonstrates a different increase in score compared to baselines. These preliminary results suggest that the incorporation of both contrastive bridging objects and language templates is crucial for designing effective Contrastive Chain-of-thought demonstrations. Additionally, setting (4), which includes synthetic examples, shows a significant increase in score. This indicates that DCD can effectively utilize automatic synthetic contrastive prompting generation with an external LLM, such as GPT-3.5.
@@
Dataset Model Greedy CP CD DoLA DCD (Ours)
Drop Quant Both
GSM8K Llama2-7B 14.32 14.25 15.39 14.03 17.28 16.00 16.00
DeepSeek-7B 12.74 14.40 - 10.37 15.47 16.38 16.38
Mistral-7B 42.23 38.90 - 43.60 48.98 47.20 48.60
Llama2-13B 29.42 25.78 32.83 28.81 33.21 31.30 32.20
StrategyQA Llama2-7B 60.04 59.91 61.62 64.02 65.15 63.18 63.32
DeepSeek-7B 60.00 59.00 - 55.10 62.40 62.01 62.01
Mistral-7B 69.04 67.73 - 70.74 74.02 72.71 73.41
Llama2-13B 65.20 66.10 69.90 68.47 71.10 70.60 70.90
Table 1: Accuracy score comparison of DCD with other existing methods: CP Chia et al. (2023), CD Li et al. (2023b), and DoLA Chuang et al. (2023). DCD variants include Drop (Dropout), Quant (Quantization), and Both (Dropout and Quantization). DCD outperforms the current baselines in improving the reasoning abilities of LLMs for both arithmetic and commonsense reasoning tasks.
@@ 5.2 Baselines
We compare DCD with three decoding intervention baselines: CP, CD, and DoLA Chuang et al. (2023). For each baseline, we adhered to the original setup’s hyperparameters. For CD, we set $\alpha$ to 0.1 and $\beta$ to 0.5. The original work of DoLA Chuang et al. (2023) only reports the setting for Llama1, so we report the best hyperparameters we can find: exit layers ranging from 00 to $14$ for 7B models and from 00 to $18$ with a step of 13B models, both with the step of 2. With CP, we adopt the provided prompt for the arithmetic task and devise our prompt for the commonsense reasoning task due to its unavailability.
@@ 5.3 Models and Hyperparameters
We conducted experiments with DCD on the Llama 1&2 Touvron et al. (2023a, b), Mistral Jiang et al. (2023), and DeepSeek DeepSeek-AI et al. (2024) models. For the Llama models, we engaged both the 7B and 13B variants. Meanwhile, we utilized the 7B versions for both Mistral and DeepSeek.
@@ 6 Results
Figure 6: Comparison of average generated token of different methods on Llama2-7B model, which clearly demonstrates DCD’s superior efficiency in achieving higher performance with less token generation compared to alternative techniques.
@@ 7.2 Quantization Amateur Model
7.3 No Distillation
To examine the impact of distillation on the amateur model, we compared our proposed DCD method with our initial experiment: a non-distilled approach combining CD with CP (Table 2). This analysis was motivated by two key hypotheses: (1) distillation applied to the amateur model could enhance the contrast between amateur and expert responses, leading to improved overall effectiveness; and (2) this improvement could match or exceed the gains achieved by utilizing a smaller amateur model, effectively removing the need for a separate smaller model in favor of the distilled amateur model.
The results of this comparison reveal that DCD consistently outperforms the non-distilled CP+CD method across various tasks. For example, when using Llama2-7B, DCD demonstrates a notable improvement over CP+CD, achieving a $1.28\%$ gain on GSM8K and a substantial $2.92\%$ gain on StrategyQA. This performance advantage is further solidified with the Llama2-13B, where DCD consistently outperforms CP+CD, achieving a $1.59\%$ increase on GSM8K and a $1.45\%$ increase on StrategyQA.
Dataset Model CP+CD DCDDrop
GSM8K Llama2-7B 16.00 17.28
Llama2-13B 31.62 33.21
StrategyQA Llama2-7B 63.23 65.15
Llama2-13B 69.65 71.10
Table 2: A comparison of accuracy scores between our initial experiment CP+CD and DCDDrop. Applying distillation significantly enhances performance over non-distilled approarches on both arithmetic and commonsense reasoning tasks.
@@ 9 Limitation and Future Work
While our study has provided valuable insights into the effectiveness of DCD, it is crucial to acknowledge certain limitations that need to be addressed.
@@ Appendix A Components of a Chain-of-thought Demonstration
Appendix B Full Prompts for Experts Model
@@ B.2 StrategyQA
Appendix C Full Prompts for Amateurs Model
@@ 42. If they ate 35, how many pieces do they
will be 3 dollars. So she has 23 - 15 dollars
left. 5 x 3 = 15. The answer is 15.
@@ C.1.2 Rule-based Number Shuffle with Calculation Error
How many trees did the grove workers plant
today?
@@
Q: Leah had 32 chocolates and her sister had
42. If they ate 35, how many pieces do they
have left in total?
A: Originally, Leah had 42 chocolates. Her
sister had 32. So in total they had 32 - 42
= -8. After eating 40, they had -8 - 40 =
40. The answer is 40.
@@ 42. If they ate 35, how many pieces do they
Q: Shawn has five toys. For Christmas, he got
two toys each from his mom and dad. How many
toys does he have now?
@@
Q: There were nine computers in the server
room. Five more computers were installed each
day, from monday to thursday. How many
computers are now in the server room?
@@
A: Olivia had 5 dollars. 23 bagels for 8
dollars each will be 5 x 8 = 39 dollars. So she
has 23 - 39 dollars left. 23 - 39 is 1. The
answer is 1.
@@ C.1.3 Rule-based Number Shuffle with Irrelerive objects and Exchange Sign
Q: There are 15 trees in the grove. Grove
workers will plant trees in the grove today.
After they are done, there will be 21 trees.
How many trees did the grove workers plant
today?
A: There were 21 apples in the basket. Later,
15 oranges were added to the basket. Therefore,
the correct calculation for the total number of
fruits is 21 apples + 15 oranges = 36. The
answer is 36.
@@
airplanes arrived. So, according to quirky
logic, 3 bicycles + 2 airplanes = 10. The
answer is 10.
Q: Leah had 32 chocolates and her sister had
42. If they ate 35, how many pieces do they
have left in total?